OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal... Please help me to simplify this Rational Expression - 8 3 ------ - ----- x^2-4 x-2

OpenStudy (abbycross167):

@jabez177 can you try to help me please?

OpenStudy (mathmale):

Abby: Factor x^2 - 4, and then determine the Lowest Common Denominator of this pair of algebraic fractions. You must modify the 2nd fraction so that its denom. becomes the LCD. Pls show your work.

OpenStudy (abbycross167):

so x^2-4 factors to (2 + x)(-2 + x)

OpenStudy (mathmale):

But (2 + x)(-2 + x) gives us 4 + 4x + x^2, which is not the same as x^2 - 4. Please try again. Note that your factors should begin with x, not with 2: 2-x is incorrect x-2 is correct

OpenStudy (jdoe0001):

well. .hmmm , if you factor \(x^2-4\implies x^2-2^2\implies ?\)

OpenStudy (jdoe0001):

so, recall \(\bf \textit{difference of squares} \\ \quad \\ (a-b)(a+b) = a^2-b^2\qquad \qquad a^2-b^2 = (a-b)(a+b)\) as mathmale said, factor that, that would help getting the LCD, so you can subtract the fractions

OpenStudy (sshayer):

(2+x)(-2+x) is same as (x+2)(x-2)

OpenStudy (abbycross167):

@jdoe0001 I'm a little confused...